# SymBoltz.jl: a symbolic-numeric, approximation-free and differentiable linear Einstein-Boltzmann solver

### v2

@@ 5 Future work
  • separation into finer computational stages (see section 2.1.3);
  • gauge transformation of equations (e.g. Newtonian $\leftrightarrow$ synchronous gauge), as already possible in CAMB [^15] https://camb.readthedocs.io/en/latest/symbolic.html;
  • perturbation theory utilities for deriving initial conditions and approximation schemes for new models;
  • transformation of differential equations with respect to conformal time $\tau$ to another (one-to-one) independent variable (e.g. to get observables as direct functions of redshift $z$);
@@
  • symbolic tensor manipulation to aid work with modified gravity theories (as a standalone package).
These tasks traditionally involve manual error-prone calculations, so such utilities could ease exploration of extended models.
We particularly aim to improve performance for differentiable runs to the level of standard runs. An attractive milestone is fast reverse-mode gradients of scalar loss functions for gradient-based MCMCs. We anticipate the symbolic structure and Jacobian generation to be useful for this, as writing $\boldsymbol{u}(\tau,\boldsymbol{p})$ in the ODE (3) and differentiating it for $\boldsymbol{v}_{i}=\partial\boldsymbol{u}/\partial p_{i}$ gives $\mathrm{d}\boldsymbol{v}_{i}/\mathrm{d}\tau=\boldsymbol{J}\boldsymbol{v}_{i}+\partial\boldsymbol{f}/\partial p_{i}$, which could be constructed analytically.
It is left for future work to generalize the code from scalar to vector and tensor perturbations, from flat to curved spacetimes, and from adiabatic to other initial conditions. Some additional models like quintessence dark energy and Brans-Dicke gravity are in the works, and we invite others to add extended models.
We hope SymBoltz can foster growth of modular interoperating packages for cosmology in Julia. For example, applications like higher-order perturbation theory (e.g. CLASS-PT by Chudaykin et al. (2020) and PyBird by D’Amico et al. (2021)), nonlinear boosting (e.g. HALOFIT by Takahashi et al. (2012) and HMCODE by Mead et al. (2021)), CMB lensing (e.g. CMBLensing.jl by Millea et al. (2020)) and nonlinear $N$-body simulations all rely on output from Boltzmann solvers. Supporting automatic differentiation across language barriers is hard, and an ecosystem of such packages in one language with support for automatic differentiation, like Julia or Python with JAX, could create a modular and differentiable cosmological modeling toolbox. A recent example of this is DISCO-DJ’s coupling of differentiable Boltzmann and $N$-body codes for end-to-end modeling from linear perturbations to non-linear structure formation (Hahn et al., 2024; List et al., 2025).
@@ 6 Conclusion
SymBoltz is a fresh Julia package for solving the linear Einstein-Boltzmann equations. It features a symbolic-numeric interface where models are built from symbolic equations and compiled to efficient numerical code. This lets modelers prototype new models with few lines code at a high level. It relaxes all approximation schemes found in older codes by solving a single set of stiff equations at all times with implicit integrators. This simplifies the equations and remains performant due to high-order implicit solvers with fast generated code for ODEs and analytical and sparse Jacobians. It is differentiable, so one can get accurate derivatives of any output quantity with respect to any input parameter. Output power spectra agree with CLASS to $0.1\%$ with standard precision and can be used fit to current data.
@@
Figure 8: A symbolic-numeric interface, approximation-freeness and differentiability are three main features of SymBoltz that form a synergy. Differentiability computes the ODE Jacobian $\boldsymbol{J}$ accurately and efficiently; which is used by implicit solvers to integrate stiff ODEs without approximations; which makes it easier to write down equations in simple symbolic form; which is used to find the sparsity pattern of $\boldsymbol{J}$; which speeds up $\boldsymbol{J}$ and the implicit solver. This is a self-reinforcing design.
Other recent codes like PyCosmo, Bolt.jl and DISCO-EB also incorporate some of these features. SymBoltz combines them and reinforces its design with a synergy between them, as explained in fig. 8. Together, these new codes offer a simpler alternative in high-level languages that complements traditional codes like CAMB and CLASS that are historically highly tuned around approximations in low-level languages. However, new codes need more work to reach the level of features and maturity of CAMB and CLASS that have grown over many years.
SymBoltz version 0.14.1 includes the flat Newtonian gauge metric, General Relativity, cold dark matter, photons, baryons and RECFAST recombination, massless and massive neutrinos, the cosmological constant and $w_{0}w_{a}$ dark energy with scalar perturbations, distances and matter and CMB spectra. Forward-mode automatic differentiation works for single-run purposes like Fisher forecasting. Fast reverse-mode scalar loss gradients, particularly through the perturbations, is a future milestone that could challenge emulators in gradient-based MCMC sampling. No differentiable Boltzmann code is yet capable of this.
SymBoltz is easy to install from https://github.com/hersle/SymBoltz.jl. Documentation [^16] https://hersle.github.io/SymBoltz.jl/ is available, and the code is tested with continuous integration to remain correct and up-to-date (see appendix C). All questions and contributions are welcome in the repository! We hope SymBoltz proves useful and that it offers valuable ideas on the Boltzmann solver market.
Acknowledgments
I thank the Julia community for creating an open ecosystem of scientific packages underlying SymBoltz. I thank Aayush Sabharwal and Christopher Rackauckas for developing ModelingToolkit.jl and OrdinaryDiffEq.jl and answering questions. I thank Hans A. Winther for helpful suggestions and feedback on the code and drafts of this paper. I thank Julien Lesgourgues, Thomas Tram and others for developing and documenting CLASS.
@@ Appendix A List of equations and practical implementation details
This appendix summarizes the equations that define the standard $\mathrm{\Lambda CDM}$ model in SymBoltz and comments on their practical implementation. We hope this can be a useful reference for others. The list is structured like SymBoltz with one component per subsection. Variables are in units where $G=c=H_{0}=1$? and temperatures are in $\mathrm{K}$, unless otherwise is stated. These units are chosen because $G$, $c$ and $H_{0}$ can be divided out from the Einstein equations as natural units, but this requires some conversion in the recombination equations that depend explicitly on $H_{0}$. In other words, times are in units of $1/H_{0}$, distances in $c/H_{0}$ and masses in $c^{3}/H_{0}G$. When $G$, $c$ or and $H_{0}$ appear explicitly in the equations, they are in SI units and used only to convert from SI units into the dimensionless units. The equations closely follow the conventions in the seminal paper by Ma and Bertschinger (1995) and very closely matches the source code of SymBoltz with Unicode characters. It is far outside the scope of this paper to derive the equations and explain the meaning of every variable.
The independent variable is conformal time $\tau$, and all derivatives $^{\prime}=\mathrm{d}/\mathrm{d}\tau$ are with respect to it (in units of $1/H_{0}$). This is perhaps the most common parametrization in the literature, and is natural because most equations are autonomous with respect to $\tau$ (i.e. $f(\tau,\boldsymbol{u})\rightarrow f(\boldsymbol{u})$, except some multipole truncation schemes that use $1/(k\tau)$, but these are unphysical). By default, integration starts from the early time $\tau=\tau_{i}=10^{-6}$, and is terminated when the scale factor crosses $a=1$, and the corresponding time $\tau=\tau_{0}$ is labeled today.
@@ A.1 Metric and spacetime $(g)$
Here $\mathscr{H}$ and $H$ are the conformal and cosmic Hubble factors (in units where $\mathscr{H}_{0}=H_{0}=1$). The scale factor $a$ is related to redshift $z$, and $\chi$ is the lookback time from today that appears in some integral solutions. SymBoltz is currently restricted to a flat spacetime.
@@ A.2 General relativity $(G)$
The gravitational theory of the $\mathrm{\Lambda CDM}$ model is General Relativity governed by the Einstein field equations $G_{\mu\nu}=8\pi T_{\mu\nu}$. By default, SymBoltz solves for the metric variables $a$, $\Phi$ and $\Psi$ with $(\mu,\nu)=(0,0)$ in the background (1st Friedmann equation) and $(\mu,\nu)=\{(0,0),(i,j)\}$ in the perturbations:
@@
Note that it is also possible to evolve other redundant combinations of the Einstein equations (such as the acceleration equation). The equations are coupled to total densities $\rho$, $\delta\rho$, pressures $P$ and anisotropic stresses $\Pi$ for an arbitrary set of species $s$ that are present in the cosmological model:
@@
We emphasize that the gravity component is completely unaware of all particle species and makes no assumptions about them. It only reacts to total stress-energy components. All species must therefore define $\rho$, $P$, $\delta\rho$, $\delta P$ and $\Pi$ explicitly, even if zero. This requirement is somewhat pedantic, but helps isolate components from each other for greater reuse when composing models.
The scale factor $a$ is initialized as the nonlinear solution of the Friedmann equation constrained to $\mathscr{H}=1/\tau$ (motivated by its radiation-dominated solution $a=\sqrt{\Omega_{r0}}\,\tau$). The constraint potential is initialized to $\Psi=20C/(15+4f_{\nu})$ with the (arbitrary) integration constant $C=1/2$ and initial energy density fraction $f_{\nu}=(\rho_{\nu}+\rho_{h})/(\rho_{\nu}+\rho_{h}+\rho_{\gamma})$ of all (massless and massive) neutrino species relative to all species that are radiative at early times. The evolved potential $\Phi$ is initialized accordingly from the constraint equation (the found solution is close to $\Phi=(1+2f_{\nu}/5)\Psi$, but providing both $\Phi$ and $\Psi$ explicitly leads to overdetermined initialization equations due to the constraint equation).
The next sections present the $\mathrm{\Lambda CDM}$ section of the ?library of species? that are available in SymBoltz.
@@ A.3 Cold dark matter $(c)$
Cold dark matter is a non-relativistic and non-interacting species that follows very simple equations:
@@
Initial conditions are adiabatic with $\delta/(1+w)=-3\Psi/2$ and $\theta=k^{2}\tau\Psi/2$. The species is parametrized by the reduced density $\Omega_{0}=\frac{8\pi}{3}\rho_{0}$ today.
@@ A.4 Baryons $(b)$
Baryons are also non-relativistic, but interact with photons through Compton scattering and are subject to recombination physics. This significantly complicates their behavior. SymBoltz currently implements equations from RECFAST [^17] https://www.astro.ubc.ca/people/scott/recfast.html version 1.5.2 (Seager et al., 1999; Wong et al., 2008; Scott and Moss, 2009):
@@
$\alpha_{\mathrm{He}_{1}}=\frac{q_{1}}{\sqrt{\frac{T_{b}}{T_{2}}}\Big(1+\sqrt{\frac{T_{b}}{T_{2}}}\Big)^{1-p_{1}}\Big(1+\sqrt{\frac{T_{b}}{T_{1}}}\Big)^{1+p_{1}}},\qquad\beta_{\mathrm{He}_{1}}=4\frac{\alpha_{\mathrm{He}_{1}}}{\lambda_{\mathrm{e}}^{3}}\exp\big({-βE_{\mathrm{He}_{1}}^{\infty,2s}}\big),\qquad K_{\mathrm{He}_{1}}=\frac{1}{K_{\mathrm{He}_{1}^{0}}^{-1}+K_{\mathrm{He}_{1}^{1}}^{-1}+K_{\mathrm{He}_{1}^{2}}^{-1}},$
@@
Every variable that does not occur on the left side of an equation is either a constant or a parameter. This includes $Y_{\mathrm{He}}$, fudge factors and wavenumbers, frequencies and energies for atomic transitions. Some important variables are the baryon temperature $T_{b}$, photon temperature $T_{\gamma}$, mean molecular weight $\mu$, baryon sound speed $c_{s}^{2}$, optical depth $\kappa$, visibility function $v$ and the free electron fraction $X_{\mathrm{e}}$ (conventionally relative to Hydrogen, so $X_{\mathrm{e}}>1$ in presence of Helium). Please consult the code and RECFAST references cited above for more details.
@@
Unlike other RECFAST implementations, SymBoltz does not approximate the stiff Peebles equations at early times by Saha approximations (although $X_{\mathrm{He}}^{++}$ is given by a Saha equation at all times). This is not necessary with a good implicit ODE solver. SymBoltz sets $C_{\mathrm{H}}=C_{\mathrm{He}_{1}}=1$ when $X_{\mathrm{e}}\gtrsim 0.99$ to avoid numerical instability at early times. Atomic calculations are done in SI units and converted to SymBoltz’ dimensionless units by factors of $H_{0}$ in SI units. The differential equation for $T_{b}^{\prime}$ is very stiff and sensitive to $T_{b}-T_{\gamma}$, but $T_{b}\approx T_{\gamma}$ in the early universe, so we rewrite it to a more stable differential equation for $\Delta T^{\prime}=T_{b}^{\prime}-T_{\gamma}^{\prime}$ instead, initialize $\Delta T=0$ and observe $T_{b}=\Delta T+T_{\gamma}$. The optical depth $\kappa(\tau)=\int_{\tau_{0}}^{\tau}\kappa^{\prime}(\tau^{\prime})\mathrm{d}\tau^{\prime}$ is really a line-of-sight integral into the past, but is integrated together with the background ODEs by initializing $\kappa(\tau_{i})=0$ to an arbitrary value, integrating the differential equation for $\kappa^{\prime}$ and subtracting the final value of $\kappa(\tau_{0})$ (i.e. $\int_{\tau_{0}}^{\tau}=\int_{\tau_{0}}^{\tau_{i}}+\int_{\tau_{i}}^{\tau}=\int_{\tau_{i}}^{\tau}-\int_{\tau_{i}}^{\tau_{0}}$). There is no tight-coupling approximation.
@@
Initial conditions are full ionization $X_{\mathrm{H}}^{+}=X_{\mathrm{He}}^{+}=1$, thermal equilibrium $T_{b}=T_{\gamma}$ ($\Delta T=0$), the arbitrary $\kappa=0$, and adiabatic perturbations $\delta/(1+w)=-3\Psi/2$ and $\theta=k^{2}\tau\Psi/2$. The baryon species is parametrized by the reduced density $\Omega_{0}=\frac{8\pi}{3}\rho_{0}$ today and the primordial Helium mass fraction $Y_{\mathrm{He}}$.
@@
SymBoltz solves thermodynamics equations together with the background equations, while some other codes treat these as separate stages. There is no meaningful performance improvement from doing this, as the size of the background (and thermodynamics) ODEs is so small. This makes a clear distinction between the background with all 0th order equations of motion, and the perturbations with all 1st order equations. It also makes it easy to create exotic models where the thermodynamics couple to the background, for example.
@@
Note that RECFAST uses fitting functions to emulate the results of more physically accurate and expensive simulations. These are tuned to work for the $\mathrm{\Lambda CDM}$ model. SymBoltz would therefore benefit from including more physically accurate recombination models for safer use with modified models.

### v3

@@ 5 Future work
  • separation into finer computational stages (see Sect. 2.1.3);
  • gauge transformation of equations (e.g., Newtonian $\leftrightarrow$ synchronous gauge), as is already possible in CAMB; [^19] https://camb.readthedocs.io/en/latest/symbolic.html
  • perturbation theory utilities to derive initial conditions and approximation schemes for new models;
  • transformation of differential equations with respect to conformal time $\tau$ to another (one-to-one) independent variable (e.g., to get observables as direct functions of redshift $z$);
@@
  • symbolic tensor algebra for work with modified gravity.
These tasks traditionally involve manual error-prone calculations, so such utilities could aid in the exploration of extended models. In particular, we also aim to improve performance for differentiable runs to the level of standard runs. An attractive milestone is fast reverse-mode gradients of scalar loss functions for gradient-based MCMCs. We anticipate the symbolic structure and Jacobian generation to be useful for this, as writing $\boldsymbol{u}(\tau,\boldsymbol{p})$ in the ODE (3) and differentiating it for $\boldsymbol{v}_{i}=\partial\boldsymbol{u}/\partial p_{i}$ gives $\mathrm{d}\boldsymbol{v}_{i}/\mathrm{d}\tau=\boldsymbol{J}\boldsymbol{v}_{i}+\partial\boldsymbol{f}/\partial p_{i}$, which could be constructed analytically.
Future work could also generalize the code from scalar to vector and tensor perturbations, from flat to curved spacetimes, and from adiabatic to other initial conditions. Some additional models for quintessence dark energy and Brans–Dicke gravity are in the works, and we invite others to add extended models.
We hope SymBoltz can foster growth of modular interoperating packages for cosmology in Julia. For example, applications such as higher order perturbation theory (e.g., CLASS-PT by Chudaykin et al. 2020 and PyBird by D’Amico et al. 2021), nonlinear boosting (e.g., HALOFIT by Takahashi et al. 2012 and HMCODE by Mead et al. 2021), CMB lensing (e.g., CMBLensing.jl by Millea et al. 2020), and nonlinear $N$-body simulations all rely on output from Boltzmann solvers. Supporting automatic differentiation across programming languages is hard, so an ecosystem of such packages in one language with support for automatic differentiation, such as Julia or Python with JAX, could create a modular and differentiable cosmological modeling toolbox. A recent example of this is DISCO-DJ’s coupling of differentiable Boltzmann and $N$-body codes for end-to-end modeling from linear perturbations to nonlinear structure formation (Hahn et al., 2024; List et al., 2025).
@@ 6 Conclusion
SymBoltz is a fresh Julia package for solving the linear Einstein–Boltzmann equations. It features a symbolic-numeric interface where models are built from symbolic equations and compiled to efficient numerical code. This lets users prototype new models with few lines of high-level code. It relaxes all approximation schemes found in older codes by solving a single set of stiff equations at all times with implicit integrators. This simplifies the equations and remains fast due to high-order implicit solvers that integrate efficient generated code for ODEs with analytical and sparse Jacobians. It is differentiable, so we can get accurate derivatives of any output quantity with respect to any input parameter. Output power spectra agree with CLASS to $0.1\%$ with standard precision, so SymBoltz can be used to fit current data.
@@
Figure 8: A symbolic-numeric interface, approximation-freeness, and differentiability are three main features of SymBoltz that form a synergy. Differentiability computes the ODE Jacobian $\boldsymbol{J}$ accurately and efficiently; which implicit solvers use to integrate stiff ODEs without approximations; which makes it easy to write equations in simple symbolic form; which provide the exact sparsity pattern of $\boldsymbol{J}$; in turn speeding up $\boldsymbol{J}$ and the solver. This loop creates a self-reinforcing design.
Other recent codes, such as PyCosmo, Bolt, and DISCO-EB, also incorporate some of these features. SymBoltz combines them and reinforces its design with a synergy between them, as explained in Fig. 8. Together, these new codes offer a simpler alternative in high-level languages that complements traditional codes such as CAMB and CLASS, which are historically highly tuned around approximations in low-level languages. However, new codes need more work to reach the level of features and maturity of CAMB and CLASS that have grown over many years.
SymBoltz version 1.0.0 includes the flat Newtonian gauge metric, general relativity, cold dark matter, photons, baryons and RECFAST recombination, massless and massive neutrinos, the cosmological constant, and $w_{0}w_{a}$ dark energy with scalar perturbations, distances, and matter and CMB spectra. Forward-mode automatic differentiation works for single-run purposes such as Fisher forecasting. Fast reverse-mode scalar loss gradients, particularly through the perturbations, is a future milestone that could challenge emulators in gradient-based MCMC sampling. No differentiable Boltzmann code is capable of this yet.
SymBoltz is publicly available [^20] https://github.com/hersle/SymBoltz.jl and easy to install. Documentation is also available, [^21] https://hersle.github.io/SymBoltz.jl/ and the code is tested with continuous integration to remain up-to-date (see Sect. C). All questions and contributions are welcome in the repository. We hope SymBoltz proves useful and that it offers valuable ideas on the Boltzmann solver market.
Acknowledgements.
I thank the Julia community for creating an extensive open ecosystem of scientific packages underlying SymBoltz. I thank Aayush Sabharwal and Christopher Rackauckas for developing ModelingToolkit and OrdinaryDiffEq and answering questions. I thank Hans A. Winther for helpful suggestions and feedback on the code and drafts of this paper. I thank Julien Lesgourgues, Thomas Tram, and others for developing and thoroughly documenting CLASS. I thank the anonymous referee for detailed and constructive comments that helped improve the quality of this paper. This research was supported by the Research Council of Norway under project number 325113.
@@ Appendix A List of equations and practical implementation details
This appendix summarizes the equations that define the standard $\mathrm{\Lambda CDM}$ model in SymBoltz and comments on their practical implementation. We hope this can be a useful reference for others. The list mirrors the internal structure of SymBoltz with one component per subsection. Unless otherwise stated, temperatures are in $\mathrm{K}$ and other variables are in units where $G=c=H_{0}=1$.? These units are chosen because $G$, $c$, and $H_{0}$ can be divided out from the Einstein equations as natural units (but this requires some conversion in the recombination equations that depend explicitly on $H_{0}$). In other words, times are in units of $1/H_{0}$, distances in $c/H_{0}$, and masses in $c^{3}/H_{0}G$. When $G$, $c$, and $H_{0}$ appear explicitly in the equations, they are in SI units and used only to convert from SI units into the dimensionless units. The equations closely follow the conventions in the seminal paper by Ma and Bertschinger (1995) and very closely match the source code of SymBoltz with Unicode characters. It is far outside the scope of this paper to derive the equations and explain the meaning of every variable.
The independent variable is conformal time $\tau$, and all derivatives $^{\prime}=\mathrm{d}/\mathrm{d}\tau$ are with respect to it (in units of $1/H_{0}$). This is perhaps the most common parameterization in the literature, and it is natural because most equations are autonomous with respect to $\tau$ (i.e., $\boldsymbol{f}(\boldsymbol{u},\tau)\rightarrow\boldsymbol{f}(\boldsymbol{u})$, except some multipole truncation schemes that have factors of $1/(k\tau)$, but these are not of physical origin). By default, integration starts from the early time $\tau=\tau_{i}=10^{-6}$ and terminates when the scale factor crosses $a=1$ at the time $\tau=\tau_{0}$ today.
@@ A.1 Metric and spacetime $(g)$
Here, $\mathscr{H}$ and $H$ are the conformal and cosmic Hubble factors (in units where $\mathscr{H}_{0}=H_{0}=1$). The scale factor $a$ is related to redshift $z$, and $\chi$ is the lookback time from today that appears in some integral solutions. SymBoltz is currently restricted to a flat spacetime.
@@ A.2 General relativity $(G)$
The gravitational theory of the $\mathrm{\Lambda CDM}$ model is general relativity governed by the Einstein field equations $G_{\mu\nu}=8\pi T_{\mu\nu}$. By default, SymBoltz solves for the metric variables $a$, $\Phi$, and $\Psi$ with $(\mu,\nu)=(0,0)$ in the background (first Friedmann equation), and $(\mu,\nu)=\{(0,0),(i,j)\}$ in the perturbations:
@@
We note that it is also possible to evolve other redundant combinations of the Einstein equations, such as the acceleration equation. The equations are coupled to total densities $\rho$ and $\delta\rho$, pressures $P$ and $\delta P$, and anisotropic stress $\Pi$ for whichever set of species $s$ that are present in the cosmological model:
@@
We emphasize that the gravity component is completely unaware of all particle species and makes no assumptions about them. It only reacts to total stress-energy components. All species must therefore define $\rho$, $P$, $\delta\rho$, $\delta P$, and $\Pi$ explicitly, even if they are zero. This requirement is somewhat pedantic, but helps isolate components from each other for greater reuse when composing models.
The scale factor $a$ is initialized as the nonlinear solution of the Friedmann equation constrained to $\mathscr{H}=1/\tau$ (motivated by its radiation-dominated solution $a=\sqrt{\Omega_{r0}}\,\tau$). The constraint potential is initialized to $\Psi=20C/(15+4f_{\nu})$ with the (arbitrary) integration constant $C=1/2$ and initial energy density fraction $f_{\nu}=(\rho_{\nu}+\rho_{h})/(\rho_{\nu}+\rho_{h}+\rho_{\gamma})$ of all (massless and massive) neutrino species relative to all species that are radiation-like at early times. The evolved potential $\Phi$ is initialized accordingly from the constraint equation (the result is close to $\Phi=(1+2f_{\nu}/5)\Psi$, but providing both $\Phi$ and $\Psi$ explicitly leads to overdetermined initialization equations that violate the constraint equation for $\Psi$). The next sections present the $\mathrm{\Lambda CDM}$ part of the ?library of species? that are available in SymBoltz.
@@ A.3 Cold dark matter $(c)$
Cold dark matter is a nonrelativistic and noninteracting species that follows very simple equations:
@@
Initial conditions are adiabatic with $\delta/(1+w)=-3\Psi/2$ and $\theta=k^{2}\tau\Psi/2$. The species is parameterized by the reduced density $\Omega_{0}=\frac{8\pi}{3}\rho_{0}$ today.
@@ A.4 Baryons $(b)$
Baryons are also nonrelativistic, but interact with photons through Compton scattering and are subject to recombination physics. This significantly complicates their behavior. SymBoltz currently implements equations from RECFAST [^22] https://www.astro.ubc.ca/people/scott/recfast.html version 1.5.2 (Seager et al. 1999; Wong et al. 2008; Scott and Moss 2009):
@@
$\alpha_{\mathrm{He}_{1}}=\frac{q_{1}}{\sqrt{\frac{T_{b}}{T_{2}}}\Big(1+\sqrt{\frac{T_{b}}{T_{2}}}\Big)^{1-p_{1}}\Big(1+\sqrt{\frac{T_{b}}{T_{1}}}\Big)^{1+p_{1}}},\qquad\beta_{\mathrm{He}_{1}}=4\frac{\alpha_{\mathrm{He}_{1}}}{\lambda_{\mathrm{e}}^{3}}\exp\big({-\beta E_{\mathrm{He}_{1}}^{\infty,2s}}\big),\qquad K_{\mathrm{He}_{1}}=\frac{1}{K_{\mathrm{He}_{1}^{0}}^{-1}+K_{\mathrm{He}_{1}^{1}}^{-1}+K_{\mathrm{He}_{1}^{2}}^{-1}},$
@@
Every variable that does not occur on the left side of an equation is either a constant or a parameter given in the code. This includes $Y_{\mathrm{He}}$, fudge factors, and wavenumbers, frequencies, and energies for atomic transitions. Some important variables defined above are the baryon temperature $T_{b}$, photon temperature $T_{\gamma}$, mean molecular weight $\mu$, baryon sound speed $c_{s}^{2}$, optical depth $\kappa$, visibility function $v$, and the free electron fraction $X_{\mathrm{e}}$ (conventionally relative to Hydrogen, so $X_{\mathrm{e}}>1$ in presence of Helium). We refer to the code and RECFAST references cited above for more details.
@@
Unlike other RECFAST implementations, SymBoltz does not approximate the stiff Peebles equations at early times by Saha approximations (although $X_{\mathrm{He}}^{++}$ is given by a Saha equation at all times). This is not necessary with a good implicit ODE solver. SymBoltz sets $C_{\mathrm{H}}=C_{\mathrm{He}_{1}}=1$ when $X_{\mathrm{e}}\gtrsim 0.99$ to avoid numerical instability at early times. Atomic calculations are done in SI units and converted to SymBoltz’ dimensionless units by factors of $H_{0}$ in SI units. The differential equation for $T_{b}^{\prime}$ is very stiff and sensitive to $T_{b}-T_{\gamma}$, but $T_{b}\approx T_{\gamma}$ in the early universe, so we rewrite it to a more stable differential equation for $\Delta T^{\prime}=T_{b}^{\prime}-T_{\gamma}^{\prime}$ instead, initialize $\Delta T=0$, and observe $T_{b}=\Delta T+T_{\gamma}$. The optical depth $\kappa(\tau)=\int_{\tau_{0}}^{\tau}\kappa^{\prime}(\tau^{\prime})\mathrm{d}\tau^{\prime}$ is really a line-of-sight integral into the past, but we solve it together with the background ODEs by initializing $\kappa(\tau_{i})=0$ to an arbitrary value, integrating the differential equation for $\kappa^{\prime}$, and subtracting the final value of $\kappa(\tau_{0})$ (i.e., $\int_{\tau_{0}}^{\tau}=\int_{\tau_{0}}^{\tau_{i}}+\int_{\tau_{i}}^{\tau}=\int_{\tau_{i}}^{\tau}-\int_{\tau_{i}}^{\tau_{0}}$). There is no tight-coupling approximation.
@@
Initial conditions are fully ionized fractions $X_{\mathrm{H}}^{+}=X_{\mathrm{He}}^{+}=1$, temperatures $T_{b}=T_{\gamma}$ ($\Delta T=0$) in equilibrium, the arbitrary $\kappa=0$, and adiabatic perturbations $\delta/(1+w)=-3\Psi/2$ and $\theta=k^{2}\tau\Psi/2$. The baryon species is parameterized by the reduced density $\Omega_{0}=\frac{8\pi}{3}\rho_{0}$ today, and the primordial Helium mass fraction $Y_{\mathrm{He}}$.
@@
SymBoltz solves thermodynamics equations together with the background equations, while some other codes treat these as separate stages. There is no meaningful performance improvement from doing this, as the size of the background (and thermodynamics) ODEs is so small. Solving these stages together creates a clear distinction between the background with all zeroth-order equations of motion, and the perturbations with all first-order equations. It also makes it possible to create exotic models where the thermodynamics affect the background, for example.
@@
We note that RECFAST uses fitting functions to emulate the results of more physically accurate and expensive simulations. These are tuned to work for the $\mathrm{\Lambda CDM}$ model. SymBoltz would therefore benefit from including more realistic recombination models for safer use with modified models.